# Cross-Space Adaptive Filter: Integrating Graph Topology and Node Attributes for Alleviating the Over-smoothing Problem

### 26 Jan 2024 v1

@@ 1. Introduction
To evaluate the effectiveness of CSF, we conduct comparative experiments with various baselines under different numbers of convolution layers. These experiments are conducted on both assortative and disassortative graphs to verify the superiority of CSF on different types of graphs. Compared to baselines, the results demonstrate that CSF not only successfully alleviates the over-smoothing problem by extracting and integrating information from both spaces but also improves the model performance on the downstream classification tasks, especially when dealing with disassortative graphs. In particular, on average, our CSF outperforms the best baseline by $+0.62$ on assortative graphs and $+10.39$ on disassortative graphs. These results highlight the importance of node attributes in assisting over-smoothing alleviation while at the same time promoting the effectiveness of deep GCNs. To sum up, we claim the following contributions.
@@ 5. Empirical Evaluation
To assess CSF’s ability to mitigate the over-smoothing issue and promote the effectiveness power of deep GCN, we conduct comparative experiments with various baselines under different numbers of convolution layers (see Section 5.2). Moreover, we reveal more characteristics of CSF by conducting ablation studies on graph topology and node attributes (see Section 5.3).
@@ 5.1. Experiment Setup
Table 1. Performance of all methods for node classification averaged over different numbers of convolution layers (i.e., $\{2,5,10,20\}$). The ”*” symbol indicates that we ignored the memory overflow situation of the corresponding model. The best results are marked in bold and the second best are underlined. For more detailed results, please refer to Table 6 in the Appendix. The results indicate CSF improves the expressiveness of deep GCN, especially when dealing with disassortative graphs.
@@
Implementation Details. The codes for most baseline methods are publicly available online. However, since there is no code available for GCN-BC, we implemented this method based on the propagation rules mentioned in their paper. Regarding the MLP, each layer is implemented by a fully connected network with ReLu activation and dropout. In the case of our proposed CSF, we utilized node attributes to construct a KNN graph with the top-20 neighbors [^7] Refer to Appendix B for the robustness analysis on the KNN graph.. To calculate edge similarity among nodes, we used a Gaussian kernel with the kernel bandwidth set to the squared average distance among nodes, without further tuning. We set $a_{2}=0.1$ for the assortative graphs, $a_{2}=100$ for the disassortative graphs, and $a_{3}=1$ for all graphs (cf. Appendix C for the analysis on $a_{2}$ and $a_{3}$). To simulate the deep GNNs, we tune the number of convolution layers of each method (i.e., $\{2,5,10,20\}$). Moreover, unlike FAGCN(Bo et al., 2021), we do not tune the hidden size on each dataset. Instead, we followed previous work (Wang et al., 2021; Kang et al., 2022; Liu et al., 2022) and set the hidden size to 16 for all methods. We also note that the learning rate in previous studies is either too large (e.g., (Li et al., 2022) directly uses $\{0.5\}$) or too small (e.g., (Bo et al., 2021) tunes the learning rate in $\{0.01,0.005\}$). Therefore, we tune the learning rate to fully evaluate the performance of all methods when the learning rate is large or small, namely $\{0.03,0.02,0.01,0.005,0.1,0.2,0.3,0.4,0.5\}$. Finally, all experiments were conducted on a machine with an Intel(R) Core(TM) i5-12400F, 16GB memory, and GeForce RTX 3060. We implemented all methods in Pytorch and optimized them using the Adam optimizer with a dropout rate of 0.5. Finally, we ran 150 epochs and selected the model with the highest validation accuracy for testing. We report the mean and variance of the results over ten runs, where all experiments were conducted under the same fixed random seed to ensure that different methods were performed with the same labeled data in each run.
@@ 5.2. Main results
Figure 5. Ablation studies on node attribute space. We report the averaged performance across various number of layers (i.e., Y-axis). Fig.(a) shows CSF effectively balances the advantages of two spaces, while Fig.(b) highlights the importance of high-frequency information in attribute space.
@@ 5.3. Ablation studies
We also analyze the parameter sensitivity and robustness of $K_{attr}$ in the Appendix C and Appendix B, respectively.
@@ Appendix E Appendix for over-smoothing problem evaluation
Table 6. Over-smoothing problem evaluation of all methods. ’OOM’ means ’out-of-memory’.

### 10 Feb 2024 v2

@@ 1. Introduction
To evaluate the effectiveness of CSF, we conduct comparative experiments with various baselines under different numbers of convolution layers. These experiments are conducted on both assortative and disassortative graphs to verify the superiority of CSF on different types of graphs. Compared to baselines, the results demonstrate that CSF not only successfully alleviates the over-smoothing problem by extracting and integrating information from both spaces but also improves the model performance on the downstream classification tasks, especially when dealing with disassortative graphs. In particular, on average, our CSF outperforms the best baseline by $+0.62$ on assortative graphs and $+10.39$ on disassortative graphs. These results highlight the importance of node attributes in assisting over-smoothing alleviation while promoting the effectiveness of deep GCNs. We claim the following contributions.
@@ 5. Empirical Evaluation
To assess CSF’s ability to mitigate the over-smoothing issue and promote the effectiveness power of deep GCN, we conduct comparative experiments with various baselines under different numbers of convolution layers (see Section 5.2). Moreover, we reveal more characteristics of CSF by conducting ablation studies on graph topology and node attributes (see Section 5.3). We also analyze the parameter sensitivity and robustness of $K_{attr}$ in the Appendix C and Appendix B, respectively.
@@ 5.1. Experiment Setup
Table 1. Performance of all methods for node classification averaged over different numbers of convolution layers (i.e., $\{2,5,10,20\}$). The ”*” symbol indicates that we ignored the memory overflow situation of the corresponding model. The best results are marked in bold and the second best are underlined. For more detailed results, please refer to Table 6 in the Appendix.
@@
Implementation Details. The codes for most baseline methods are publicly available online. Regarding the MLP, each layer is implemented by a fully connected network with ReLu activation and dropout. In the case of our proposed CSF, we utilized node attributes to construct a KNN graph with the top-20 neighbors [^7] Refer to Appendix B for the robustness analysis on the KNN graph.. To calculate edge similarity among nodes, we used a Gaussian kernel with the kernel bandwidth set to the squared average distance among nodes, without further tuning. We set $a_{2}=0.1$ for the assortative graphs, $a_{2}=100$ for the disassortative graphs, and $a_{3}=1$ for all graphs (cf. Appendix C for the analysis on $a_{2}$ and $a_{3}$). To simulate the deep GNNs, we tune the number of convolution layers of each method (i.e., $\{2,5,10,20\}$). Moreover, unlike FAGCN(Bo et al., 2021), we do not tune the hidden size on each dataset. Instead, we followed previous work (Wang et al., 2021; Kang et al., 2022; Liu et al., 2022) and set the hidden size to 16 for all methods. We also note that the learning rate in previous studies is either too large (e.g., (Li et al., 2022) directly uses $\{0.5\}$) or too small (e.g., (Bo et al., 2021) tunes the learning rate in $\{0.01,0.005\}$). Therefore, we tune the learning rate to fully evaluate the performance of all methods when the learning rate is large or small, namely $\{0.03,0.02,0.01,0.005,0.1,0.2,0.3,0.4,0.5\}$. Finally, all experiments were conducted on a machine with an Intel(R) Core(TM) i5-12400F, 16GB memory, and GeForce RTX 3060. We implemented all methods in Pytorch and optimized them using the Adam optimizer with a dropout rate of 0.5. Finally, we ran 150 epochs and selected the model with the highest validation accuracy for testing. We report the mean and variance of the results over ten runs, where all experiments were conducted under the same fixed random seed to ensure that different methods were performed with the same labeled data in each run.
@@ 5.2. Main results
Figure 5. Ablation studies on node attribute space. We report the averaged performance across various number of layers.
@@ Appendix E Appendix for over-smoothing problem evaluation
Table 6. Over-smoothing problem evaluation of all methods. ’OOM’ means ’out-of-memory’. GPRGNN (Chien et al., 2020) and ChebNetII (He et al., 2022) are tailored for disassortative graghs.